# Perm: A Parametric Representation for Multi-Style 3D Hair Modeling

### v4

@@ D.3 Single-view Hair Reconstruction
Figure 31: Single-view hair reconstruction results of HAAR (Sklyarova et al., 2023b).
@@ D.4 Hair-conditioned Image Generation
In Fig. 32 we present a comparison of image generation results with and without our input hair conditions. Note that we use the pre-visualization from MeshLab (Cignoni et al., 2008) as the structural image condition, which we found to perform better than the final renderings. Without our hair conditions, images generated with rough text prompts like “wavy and short hair” cannot guarantee the production of the desired hairstyle, and their hairstyles often vary with pose, resulting in a loss of multi-view consistency. Leveraging the rich information encoded in the 3D hair geometry, our hair conditions effectively address these issues, producing high-quality portrait images with a more consistent hairstyle.
@@
In Fig. 33 we further showcase some generated images conditioned on various input hairstyles. The texture colors in the input hairstyle renderings are only for aesthetic purpose and have no effect on the skin or hair colors in the generated images. Although structural conditioned image generation is not our focus and needs more future investigation on improving the structural alignment, this application reveals the potential of deploying current T2I models as a “neural renderer” (Tewari et al., 2020) within the traditional CG pipeline.
@@
Input Hairstyles Generated by Perm and Simulated by (Amador Herrera et al., 2024)

### v5

@@ D.3 Single-view Hair Reconstruction
Figure 31: Single-view hair reconstruction of HAAR (Sklyarova et al., 2023b).
@@ D.4 Hair-conditioned Image Generation
In Fig. 32 we present a comparison of image generation results with and without our input hair conditions. Note that we use the pre-visualization from MeshLab (Cignoni et al., 2008) as the structural image condition, which we found to perform better than the final renderings. Without our hair conditions, images generated with rough text prompts like “wavy and short hair” cannot guarantee the production of the desired hairstyle, and their hairstyles often vary with pose, resulting in a loss of multi-view consistency. Leveraging the rich information encoded in the 3D hair geometry, our hair conditions effectively address these issues, producing high-quality portrait images with more consistent hairstyles.
@@
In Fig. 33 we further showcase some generated images conditioned on various input hairstyles. The texture colors in the input hairstyle renderings are only for aesthetic purposes and have no effect on the skin or hair colors in the generated images. Although structural conditioned image generation is not our focus and needs more future investigation on improving the structural alignment, this application reveals the potential of deploying current T2I models as a “neural renderer” (Tewari et al., 2020) within the traditional CG pipeline.
@@
Input Hairstyles Generated by Perm and Simulated by (Herrera et al., 2024)